1. There are three brands, say X, Y and Z of an item available in the market. A consumer chooses exactly one of them for his use. He never buys two or more brands simultaneously. The probabilities that he buys brands X, Y and Z are 0.20, 0.16 and 0.45, respectively.

a) What is the probability that he does not buy any of the brands?

b) Given that a customer buys some brand, what is the probability that he buys brand


1
Expert's answer
2021-06-24T12:38:39-0400

Given


"P(X\cap Y)=P(X\cap Z)=P(Y\cap Z)=P(X\cap Y\cap Z)=0"

Then


"P(X\cup Y\cup Z)=P(X)+P(Y)+P(Z)"

"-P(X\cap Y)-P(X\cap Z)-P(Y\cap Z)"

"+P(X\cap Y\cap Z)"

"=P(X)+P(Y)+P(Z)-0-0-0+0"

"=P(X)+P(Y)+P(Z)"

a)


"P(X'\cap Y'\cap Z')=1-P(X\cup Y\cup Z)"

"=1-(P(X)+P(Y)+P(Z))"

"=1-(0.20+0.16+0.45)=0.19"


b)

"P(X|(X\cup Y\cup Z))=\dfrac{P(X\cap(X\cup Y\cup Z))}{P(X\cup Y\cup Z)}"

"=\dfrac{P(X)}{P(X)+P(Y)+P(Z)}"

"=\dfrac{0.20}{0.20+0.16+0.45} =\dfrac{20}{81}\approx0.2469"